Question 6


Let u =< −2,1,−1,v =< −3,2,−1 >and w =< 1,3,5 >. Compute:


(6.1) u ×w,


(6.2)u ×(w ×v)and(u × w)×v.


1
Expert's answer
2021-06-29T12:30:19-0400

6.1) "u×w=\begin{vmatrix}
 i & j &k \\
 -2 & 1 & -1\\
1 & 3 & 5

\end{vmatrix}" "=i(5+3)-j(-10+1)+k(-6-1)\n=8i+9j-7k" "6.2) u\u00d7(w\u00d7v)=(u.v)w-(u.w)v\n={<-2,1,-1>.<-3,2,-1>}<1,3,5>-{<-2,1,-1>.<1,3,5>}<-3,2,-1>\n=(6+2+1)<1,3,5>-(-2+3-5)<-3,2,-1>\n=<9,27,45>-<12,-8,4>=<-3,35,41>" "\to (u×w)×v=<8,9,-7>×<-3,2,-1>=\begin{vmatrix}
 i & j &k \\
 8 & 9 & -7\\
-3 & 2 & -1
\end{vmatrix}""=i(-9+14)-j(-8-21)+k(16+27)\n=5i+29j+43k"